OpenStudy (anonymous):

Can you simplify sin^2x to x(sin (x) )

myininaya (myininaya):

no

myininaya (myininaya):

\[\sin^2(x)=\sin(x)\cdot \sin(x)\neq x \cdot \sin(x)\]

OpenStudy (anonymous):

it can be rewritten as 1 - cos^2x and derivated to 2sinxcosx... but not xsinx

OpenStudy (anonymous):

sorry i meant to say 

myininaya (myininaya):

\[\sin(x) \neq x \]

OpenStudy (anonymous):

ok got it, so then how would you simplify it?

myininaya (myininaya):

how can you simplify it?

myininaya (myininaya):

it looks pretty simplified to me

OpenStudy (anonymous):